### v1

# FakeNewsGPT4: Advancing Multimodal Fake News Detection through Knowledge-Augmented LVLMs

Abstract
The massive generation of multimodal fake news exhibits substantial distribution discrepancies, prompting the need for generalized detectors. However, the insulated nature of training within specific domains restricts the capability of classical detectors to obtain open-world facts. In this paper, we propose FakeNewsGPT4, a novel framework that augments Large Vision-Language Models (LVLMs) with forgery-specific knowledge for manipulation reasoning while inheriting extensive world knowledge as complementary. Knowledge augmentation in FakeNewsGPT4 involves acquiring two types of forgery-specific knowledge, i.e., semantic correlation and artifact trace, and merging them into LVLMs. Specifically, we design a multi-level cross-modal reasoning module that establishes interactions across modalities for extracting semantic correlations. Concurrently, a dual-branch fine-grained verification module is presented to comprehend localized details to encode artifact traces. The generated knowledge is translated into refined embeddings compatible with LVLMs. We also incorporate candidate answer heuristics and soft prompts to enhance input informativeness. Extensive experiments on the public benchmark demonstrate that FakeNewsGPT4 achieves superior cross-domain performance compared to previous methods. Code will be available.
FakeNewsGPT4: Advancing Multimodal Fake News Detection through Knowledge-Augmented LVLMs
[…]
1 Introduction
The wide spread of fake news has become an important social issue, posing threats to politics Fisher et al. (2016), finance ElBoghdady (2013), and public health Naeem and Bhatti (2020). Misusing advanced generative models to create misinformation further exacerbates these issues, manifested in the rise of visual deepfakes Zhou et al. (2021) and the proliferation of text fake news Sheng et al. (2021). Furthermore, fake news through multimedia convergence disseminates more expansive information with greater impact to mislead readers. Thus, it is of utmost urgency to develop methods for multimodal fake news detection (MFND).
Figure 1: Illustration of leveraging world knowledge from LVLMs to mitigate the issue of domain shift. (a) Distribution discrepancies of fake news manifested in both manipulated content and realistic context. (b) Profound understanding of unseen fake news aided by the world facts from the LVLMs’ knowledge bases.
Current MFND methods Ying et al. (2023); Chen et al. (2023); Shao et al. (2023a) primarily confine training to given domains (i.e., closed systems), leading to a limited acquisition of information. However, open-world fake news exhibits substantial distribution discrepancies Zhu et al. (2022b), termed domain shift Nan et al. (2021); Zhu et al. (2022c), which is manifested in two aspects in Fig. 1-(a). 1) Diverse manipulated content. Any part of news can be manipulated through creative strategies driven by anonymous intentions. These strategies can selectively distort emotional tendencies, reported incidents, key entities, etc. 2) Diverse realistic context. The real-world context of news differs in regional perspectives and thematic focus. For instance, the common words with distinct styles in BBC are “Prime Minister” and “Queen”, in contrast to “President” and “Obama” in the Washington Post. Hence, it is challenging for existing models in the closed systems to understand and characterize open-world fake news.
Considering the above challenges, we propose to leverage world knowledge from Large Vision-Language Models (LVLMs) Liu et al. (2023); Zhu et al. (2023) to provide complementary information for fake news detectors. This knowledge encompasses a wide array of world facts Bubeck et al. (2023) about public celebrities, social events etc, which enables a comprehensive understanding of unseen fake news. As depicted in Fig 1-(b), the agnostic instance "Donald Trump" can be precisely identified for robust semantic analysis aided by inherent world knowledge from LVLMs.
Despite being proficient in recognizing common instances, LVLMs lack forgery-specific knowledge Zhao et al. (2023), compromising their effectiveness in MFND tasks (see Fig. 2-(a)). Two types of external information are crucial for manipulation reasoning and should be incorporated into the training and inference of vision-language models Zhu et al. (2022a): semantic correlation and artifact trace. On the one hand, manipulated media collectively alter cross-modal correlation, leading to semantic inconsistency between modalities Shao et al. (2023a). On the other hand, manipulated images typically leave suspicious artifact traces Shiohara and Yamasaki (2022).
Figure 2: Illustration of the effect of forgery-knowledge augmentation. (a) Vanilla LVLMs struggle to correctly judge the news veracity. (b) Incorporating external knowledge (i.e., semantic correlation and artifact trace) into LVLMs helps models make accurate predictions.
To enhance cross-domain performance for multimodal fake news detection, in this paper, we present a novel framework, namely FakeNewsGPT4 which augments LVLMs with forgery-specific knowledge. This framework incorporates external information that is beneficial for manipulation reasoning and capitalizes on the world knowledge from LVLMs as complementary. As shown in Fig. 2-(b), we first establish representations of forgery-specific knowledge by cross-modal reasoning and fine-grained verification. Specifically, the multi-level cross-modal reasoning module utilizes cross-modal interactions to focus on regions relevant to semantic correlations. Additionally, the dual-branch fine-grained verification module gathers local spatial information to encode artifact traces through supervised guidance. Subsequently, the generated knowledge is transformed into refined embeddings to align with the language space of LVLMs. We devise MFND instruction-following data for fine-tuning and employ both candidate answer heuristics and soft prompts to unleash the extensive knowledge of language models.
Our contributions are summarized as follows:
  • We pioneer leveraging world knowledge from large vision-language models (LVLMs) to tackle the domain shift issue in multimodal fake news detection. Building upon capitalizing on the knowledge from LVLMs, we propose a generalized detector, FakeNewsGPT4.
  • FakeNewsGPT4 augments LVLMs with forgery-specific knowledge for manipulation reasoning. We propose cross-modal reasoning and fine-grained verification to extract both semantic correlations and artifact traces. To merge the generated knowledge into LVLMs, we also devise MFND instruction data along with answer heuristics and soft prompts.
  • The extensive experiments demonstrate the effectiveness of our proposed method under multiple cross-domain settings.
2 Related Work
Figure 3: Architecture of our proposed FakeNewsGPT4. Given a manipulated image-text pair, the multi-level cross-modal reasoning module (a) first extracts cross-modal semantic embeddings and visual patch features. Then, these visual patch features are processed by the dual-branch fine-grained verification module (b) to encode discriminative artifact embeddings. Finally, the semantic and artifact embeddings are incorporated into the forgery-aware vision-language model (c) combined with candidate answer heuristics and soft prompts for deep manipulation reasoning.
Fake News Detection.
Fake news detection works can be categorized into unimodal (image-based and text-based) and multimodal methods. Image-based method Cao et al. (2020) proposes to exploit edited traces to verify the truth of visual content. One group of methods focuses on the spatial domain to mine artifact traces, such as blending Shiohara and Yamasaki (2022), patch consistency Zhao et al. (2021), and reconstruction Liang et al. (2022). The other works transform images into the frequency domain by applying DCT Qian et al. (2020) and phase spectrum Liu et al. (2021b).
Text-based methods primarily delve into various aspects such as sentiment Ghanem et al. (2021), user feedback Min et al. (2022), news environment Sheng et al. (2022), propaganda techniques Huang et al. (2023) and temporal patterns Hu et al. (2023). Recent studies have emphasized the importance of multi-domain Nan et al. (2021); Zhu et al. (2022c) and cross-domain detection methods Nan et al. (2022); Lin et al. (2022) in alleviating the domain shift issue.
In contrast to unimodal methods, multimodal methods adhere to incorporate cross-modal features to extract semantic representations by graph convolution Wang et al. (2020), enhanced entity Qi et al. (2021), co-attention layers Wu et al. (2021), contextual attention Qian et al. (2021), ambiguity learning Chen et al. (2022), multi-view Ying et al. (2023), and causal reasoning Chen et al. (2023). A recent work Shao et al. (2023a) presents HAMMER, a powerful model that combines contrastive learning and cross-modal aggregation. However, all these methods are typically trained and deployed within closed systems, overlooking the potential benefits of accessing world knowledge.
Large Vision-Language Models.
Large language models (LLMs) Touvron et al. (2023); Chiang et al. (2023) have showcased remarkable performance on various linguistic tasks. More recently, researchers are exploring extending the capability of LLMs to perceive visual signals. LLaVA Liu et al. (2023) and Mini-GPT4 Zhu et al. (2023) first facilitate image-text feature alignment followed by visual instruction tuning. PandaGPT Su et al. (2023) employs a simple linear layer as a bridge between ImageBind Girdhar et al. (2023) and the Vicuna model Chiang et al. (2023), allowing for the multimodal input.
Knowledge Augmented Language Models.
Utilizing external knowledge to augment language models (LMs) has emerged as a promising solution in knowledge-intensive tasks Kang et al. (2023). One line of works is retrieval-augmented LMs which retrieve relevant passages and incorporate them into LMs using chunked cross-attention Borgeaud et al. (2022), dual instruction tuning Lin et al. (2024) and self-reflection Asai et al. (2024). Nevertheless, the potential of LVLMs augmented with forgery-specific knowledge in multimodal fake news detection remains unexplored.
3 Proposed Method
In this section, we present a unified framework named FakeNewsGPT4 which incorporates forgery-specific knowledge into LVLMs for manipulation reasoning. We first introduce the overall framework architecture. Then, we elaborate on our proposed multi-level cross-modal reasoning module, dual-branch fine-grained verification module, and forgery-aware vision-language model. Finally, we detail the loss function for training.
3.1 The Overall Framework
The comprehensive architecture of FakeNewsGPT4 is illustrated in Fig. 3. Given a manipulated image-text pair $\left(I,T\right)$, we design two separate modules to extract semantic correlations and artifact traces as forgery-specific knowledge representations. Specifically, the multi-level cross-modal reasoning module utilizes dual-branch cross-attention mechanisms to guide cross-modal interactions, facilitating the encoding of semantic embeddings. Concurrently, the dual-branch fine-grained verification module collects local spatial information to establish artifact embeddings through supervised verification. Then, the forgery-aware vision-language model leverages both the forgery-specific knowledge and inherent world knowledge for deep manipulation reasoning.
3.1.1 Multi-level Cross-modal Reasoning
The input image and text are encoded first by the frozen pre-trained encoders, denoted as $E_{v}$ for the image and $E_{t}$ for the text. To obtain both low-level elements and high-level semantic cues, we partition the image encoder into multiple layers, enabling the integration of intermediate patch-level features. As shown in Fig. 3-(a), the image features $f_{v}$ and text features $f_{t}$ are separately represented as:
$f_{v}=E_{v}\left(I\right),\,\,f_{t}=E_{t}\left(T\right).$ (1)
Both features contain information of object instances within a single modality but lack prior insight into objects referenced by the other modality. This absence of complementary information between modalities may hinder cross-modal semantic reasoning. To this end, we devise the dual-branch cross-attention to guide the interaction between visual and textual features, enabling the extraction of semantic correlations. Attention function Vaswani et al. (2017) is performed on normalized query ($Q$), key ($K$), and value ($V$) features as:
$\begin{split}\text{Attention}(Q,K,V)=\text{Softmax}(\frac{K^{T}Q}{\sqrt{D}})V.\end{split}$ (2)
In dual-branch cross-attention, each modal feature (e.g., image) is allowed to serve as queries $Q$, while keys $K$ and values $V$ can be taken from the other modality (e.g., text):
$u_{v}=\text{Attention}\left(f_{v},f_{t},f_{t}\right),$ (3)
$u_{t}=\text{Attention}\left(f_{t},f_{v},f_{v}\right),$ (4)
where $u_{v}=\{u_{v}^{\rm cls},u_{v}^{\rm pat}\}$, $u_{t}=\{u_{t}^{\rm cls},u_{t}^{\rm pat}\}$. Here, $u_{v}^{\rm cls}$ and $u_{t}^{\rm cls}$ are [CLS] tokens from visual/textual embeddings interacted with text/image information. $u_{v}^{\rm pat}$ and $u_{t}^{\rm pat}$ are corresponding patch embeddings.
Based on the cross-modal interaction described above, the [CLS] tokens $u_{v}^{\rm cls}$ and $u_{t}^{\rm cls}$ can naturally focus on the inter-modal semantic correlations. We concatenate these two [CLS] tokens $\{u_{v}^{\rm cls},u_{t}^{\rm cls}\}$ as a joint representation of semantic embedding. Then we use a learnable linear layer to connect this generated knowledge into the word embedding space of LVLMs, facilitating profound semantic reasoning aided by world factual knowledge.
3.1.2 Dual-branch Fine-grained Verification
In addition to extracting semantic correlations between visual and textual features, mining artifact traces is also crucial for distinguishing fake news. Unlike [CLS] token, visual patch tokens with position encoding Vaswani et al. (2017) contain richer local spatial information. Given visual patch tokens $u_{v}^{\rm pat}$, we propose a dual-branch fine-grained verification module to encode them into artifact embeddings, guided by grounding annotations.
As depicted in Fig. 3-(b), the upper branch which comprises a text encoder and a pixel decoder, is designed to utilize language-driven contrastive learning for pixel-wise verification. Dense features with good language alignment can provide complementary benefits for fine-grained segmentation Jeong et al. (2023); Gu et al. (2023). Specifically, two class prompts are initially encoded by the text encoder to obtain corresponding features $F_{p}^{i}\left(i=1,2\right)$, representing “natural” and “unnatura” states. Detailed contrastive class prompts are presented in the Appendix A. To restore local spatial details, the pixel decoder with consecutive deconvolution layers converts low-resolution features $u_{v}^{\rm pat}$ into high-resolution features $F_{h}$. Since patch-level features are not aligned with the textual space, we project both textual features $F_{p}$ and visual features $F_{h}$ into a shared representation space. Subsequently, the projected features $\tilde{F}_{p}$ and $\tilde{F}_{h}$ are employed to compute similarity scores followed by upsampling:
$w^{i}=\text{Upsample}\left(\tilde{F}_{p}^{T}\tilde{F}_{h}\right),\,\,\left(i=1,2\right).$ (5)
By applying the scores spatially, we can establish the manipulated segmentation map to achieve per-pixel prediction:
$M_{s}=\sum_{i=1}^{2}{\log\left(\frac{e^{w^{i}}}{\sum{e^{w^{i}}}}\right)}.$ (6)
To enrich the representation of artifact traces, the lower branch employs multi-head attention for patch-level verification. We utilize a learnable token $q_{\mathrm{tok}}$ serving as a query, while visual features $u_{v}^{\rm pat}$ act as both the key and value. Through the multi-head attention mechanism, local information related to artifacts is aggregated within $u_{\rm agg}$ under the supervision of bounding box grounding:
$u_{{\rm agg}}=\mathrm{Attention}\left(q_{{\rm tok}},\tilde{u}_{v}^{\rm pat},\tilde{u}_{v}^{\rm pat}\right).$ (7)
To leverage the artifact knowledge contained in the two-branch features (i.e., the manipulated segmentation map $M_{s}$ and the aggregated token $u_{{\rm agg}}$), we separately devise a pixel encoder with multiple convolution layers and a simple linear layer. In this manner, both $M_{s}$ and $u_{{\rm agg}}$ are converted into artifact embeddings aligned with the final vision-language model.
3.1.3 Forgery-aware Vision Language Model
With the extraction of two types of forgery-specific knowledge, merging this external knowledge into LVLMs becomes imperative. Both the obtained semantic embeddings and artifact embeddings have been refined to lie in the embedding space for compatibility with language models. Moreover, the introduction of MFND instruction data along with answer heuristics and soft prompts can further activate the capacity of LVLMs.
Due to the lack of instruction-follow data in the MFND task, we carefully design an instruction template following the conversational format of the Vicuna Chiang et al. (2023), as shown below:
###Human: <Img><ImageFeature></Img>
<ForgeryFeature>[Human Prompt] ###Assistant:
In this prompt, <ImageFeature> represents the visual tokens produced by the image encoder and <ForgeryFeature> is the combination of semantic and artifact embeddings. The human prompt adopts the multiple choice question answering format, as shown in the dialog box of Fig. 3-(c).
To unleash the potential knowledge of LVLMs in solving MFND tasks, we devise two prompt strategies to serve as more informative inputs in Fig. 3-(c). On the one hand, we utilize candidate answer heuristics Robinson and Wingate (2022); Shao et al. (2023b) to present both the question and answer options to LVLMs and make them predict the symbol (e.g., “A”) associated with the selected answer. This approach enables the language models to explicitly compare different candidate answers showcasing more accurate responses. On the other hand, we implement soft prompt tuning to introduce learnable continuous vectors while freezing the language models. These vectors combined with semantic embeddings facilitate the extraction of additional semantic information. Meanwhile, this approach reduces the burden of LVLMs to learn forgery-aware alignment, thereby mitigating the catastrophic forgetting problem.
3.1.4 Loss Function
Two groups of loss functions are employed in our training procedure: cross-entropy loss, and two-branch verification losses.
Cross-entropy Loss.
In the training of language models, cross-entropy loss is employed to measure the disparity between the text sequence predicted by the models and the target text sequence. The formula is computed according to:
$\mathcal{L}_{\mathrm{ce}}=-\sum_{i=1}^{n}{y_{i}{\mathrm{log}}(p_{i})},$ (8)
where $n$ denotes the token count, $y_{i}$ is the true label for token $i$ and $p_{i}$ is the corresponding predicted probability.
Dual-branch Verification Loss.
Two-branch verification losses are designed for the precise encoding of artifact traces guided separately by pixel-wise and patch-level verification. Pixel-wise verification introduces focal loss Lin et al. (2017) and dice loss Milletari et al. (2016) to facilitate supervision for the manipulated segmentation map $M_{s}$: $\mathcal{L}_{\mathrm{pixel}}=\mathcal{L}_{\mathrm{focal}}+\mathcal{L}_{\mathrm{dice}}$. Patch-level verification involves regressing the final bounding box by utilizing the aggregated token $u_{{\rm agg}}$ and computing the regression losses with the ground-truth box by introducing L1 loss and GIoU loss Rezatofighi et al. (2019): $\mathcal{L}_{\text{patch}}=\mathcal{L}_{\mathrm{L}_{1}}+\mathcal{L}_{\text{giou}}$. More details about verification losses are provided in Appendix B.
Finally, the overall loss function is defined as:
$\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{pixel}}+\mathcal{L}_{\mathrm{patch}}.$ (9)
4 Experiments
In this section, we first introduce the overall experimental setup and then provide comprehensive experimental results to demonstrate the superiority of our proposed method.
Domain BBC Guardian USA Wash.
Train # Real 20436 55459 15472 12725
# Fake 20375 54244 16339 13134
Total 40811 109703 31811 25859
Test # Real 3156 9109 2533 2078
# Fake 6214 17919 5393 4303
Total 9370 27028 7926 6381
[…]
Train Method Test
BBC Guardian USA Wash.
[…]
PandaGPT 49.99 50.06 66.31 49.58 50.19 66.30 49.47 50.37 68.04 49.51 50.47 67.43
BBC PandaGPT+SPT 54.93 47.01 48.29 53.89 47.23 54.33 51.19 49.72 55.96 52.43 48.07 54.91
HAMMER 87.35 21.40 79.98 80.82 26.81 76.46 65.16 39.40 69.02 67.01 37.92 68.28
FakeNewsGPT4 89.61 18.61 81.55 84.95 23.55 77.08 73.10 33.91 70.50 74.81 32.29 71.52
Guardian PandaGPT+SPT 54.66 46.91 46.04 56.57 45.21 50.37 55.99 45.85 50.41 55.58 46.11 50.15
HAMMER 73.74 31.53 66.74 93.90 13.38 87.45 65.34 39.61 69.23 63.24 41.17 68.50
FakeNewsGPT4 82.65 25.11 74.92 93.93 13.38 86.60 74.32 32.65 71.06 73.15 33.13 70.16
USA PandaGPT+SPT 50.01 50.31 59.12 52.88 47.94 59.49 56.50 45.29 58.98 53.89 47.76 60.27
HAMMER 68.44 35.32 69.81 74.71 30.46 74.35 85.11 22.68 79.08 81.60 25.17 76.90
FakeNewsGPT4 74.17 31.23 72.91 78.82 27.63 76.66 89.64 18.69 80.96 87.76 20.25 80.68
Wash. PandaGPT+SPT 54.40 46.81 44.42 53.08 47.87 48.18 52.38 48.39 49.21 53.93 47.09 56.43
HAMMER 71.29 33.54 70.59 76.78 29.40 74.21 82.11 25.66 77.35 83.30 24.26 77.64
FakeNewsGPT4 78.56 28.52 73.47 81.97 25.31 76.29 87.07 21.19 79.06 87.94 19.81 80.16
[…]
4.1 Experimental Setup
Dataset.
[…]
Evaluation Metrics.
We treat the multimodal fake news detection problem as a binary classification task. Following previous works Shao et al. (2023a); Nan et al. (2022), we apply the Area Under the Receiver Operating Characteristic Curve (AUC), Equal Error Rate (EER), the Accuracy Score (ACC) as our evaluation metrics.
Baselines.
The proposed FakeNewsGPT4 is compared with the following strong baseline models: 1) PandaGPT Su et al. (2023): The vanilla PandaGPT model effectively aligns visual features with the text space of LLMs, enabling it to perform complex multimodal tasks in a zero-shot manner. 2) PandaGPT+SPT: This model integrates PandaGPT with soft prompt tuning Lester et al. (2021) by using the predefined question prompt along with learnable continuous vectors to fine-tune the LVLM during the instruction tuning phase. 3) HAMMER Shao et al. (2023a): HAMMER employs two unimodal encoders to encode image and text embeddings with contrastive learning for alignment, and then summarize multimodal information through the multimodal aggregator.
Method BBC & Guardian USA & Wash.
USA Wash. BBC Guardian
[…]
PandaGPT+SPT 50.59 49.56 54.44 51.95 48.68 52.55 52.38 48.09 52.54 51.74 48.75 53.13
HAMMER 63.45 39.81 69.19 62.59 40.34 68.64 73.95 32.48 71.86 80.04 27.54 76.06
FakeNewsGPT4 75.17 33.35 71.42 75.15 32.85 70.52 81.06 26.21 75.26 85.88 22.41 78.87
[…]
Implementation Details.
We utilize the visual encoder and text encoder sourced from ImageBind-Huge Girdhar et al. (2023) as the backbone for our image and text feature extractors. Moreover, we employ Vicuna-7B Chiang et al. (2023) as the inferential LLM, connected through a linear layer. The model is initialized from the instruction-tuned checkpoint provided by PandaGPT. All training images are resized to 224 × 224 and subjected to random horizontal flipping, as well as random perturbation techniques such as JPEG compression and Gaussian Blur following Shao et al. (2023a). The multi-level cross-modal reasoning module extracts intermediate patch features from the 8th, 16th, 24th, and 32nd layers of the image encoder. We set the learning rate as 1.5e-5 with a batch size of 16 and a maximum of 10 epochs when trained on the BBC subset. Linear warm-up and the one-cycle cosine learning schedule are adopted. All experiments are conducted on four NVIDIA GeForce 3090 GPUs with PyTorch. More details about hyperparameter settings are provided in Appendix D.
4.2 Performance Comparison
We evaluate the cross-domain performance of our FakeNewsGPT4 with baselines in both single-domain and multiple-domain settings respectively.
Single-domain Setting.
Table 2 presents the performance of our method and other baseline models in the challenging scenario where a single domain is available. We randomly select one subset as the source domain for training and the remaining subsets as the target domains for testing. From the results, we make the following observations: 1) The performance of the existing MFND method drops significantly when tested on the unknown subsets, which verifies the existence of domain shift caused by the deviation in propagation contents. 2) FakeNewsGPT4 yields substantial improvement on recent LVLMs, PandaGPT, and PandaGPT with soft prompt tuning, in both intra-domain and cross-domain testing. Such huge improvement demonstrates the effectiveness of forgery-specific knowledge augmentation in our framework. 3) Compared with the state-of-the-art method, HAMMER, our approach shows more remarkable improvement in cross-domain testing. For instance, for models trained on the BBC subset, FakeNewsGPT4 achieves a 7.7% increase in AUC when testing on the Washington Post subset. This may be credited to the effective utilization of inherent world knowledge from LVLMs in mitigating distribution discrepancies. The combination of forgery-specific knowledge and world knowledge facilitates profound manipulation reasoning in FakeNewsGPT4.
Multiple-domain Setting.
The inclusion of domestic news such as the BBC and The Guardian from British, as well as USA TODAY and The Washington Post from America, increases dataset diversity in practical scenarios. We select two subsets from identical countries for training and the remaining two for testing. The results are summarized in Table 3. Our FakeNewsGPT4 exhibits significant superiority over both PandaGPT using soft prompt tuning and HAMMER by a large margin. This reveals the effectiveness of our framework in instance-wise domain generalization guided by world knowledge derived from LVLMs, even when jointly learning multiple source domains.
Componet Module Test
ML&CM Reasoning DB&FG Verification Guardian USA Wash. Avg.
[…]
53.89 47.23 54.33 51.19 49.72 55.96 52.43 48.07 54.91 52.50 48.34 55.07
$\checkmark$ 83.53 23.95 77.03 67.64 36.45 69.52 70.97 34.60 70.33 74.05 31.67 72.29
$\checkmark$ 81.79 26.76 73.98 66.59 39.02 65.96 69.60 36.67 67.25 72.66 34.15 69.06
$\checkmark$ $\checkmark$ 84.95 23.55 77.08 73.10 33.91 70.50 74.81 32.29 71.52 77.62 29.92 73.03
FakeNewsGPT4 w/o ML 81.27 26.70 75.49 65.25 39.71 69.20 68.97 36.54 69.21 71.83 34.32 71.30
w/o DB (upper) 83.70 23.92 77.29 62.85 40.05 70.00 67.73 36.27 70.49 71.43 33.41 72.59
w/o DB (lower) 84.12 23.88 76.74 70.37 36.36 69.66 73.50 32.87 70.30 76.00 31.04 72.23
Table 4: Ablation study of component modules. We evaluate the AUC (in %), EER (in %), and ACC (in %) of variant models on the remaining three subsets when trained on the BBC subset. ML: the extraction of multi-level patch features in the cross-modal (CM) reasoning module. DB: the extraction of dual-branch artifact features in the fine-grained (FG) verification module. Avg. denotes the mean value on the three testing subsets.
Strategy Guardian USA Wash. Avg.
w/o SPT 84.76 71.98 74.29 77.01
w/o CAH 83.14 67.38 70.55 73.69
w/o SPT & CAH 83.32 66.12 69.94 73.13
FakeNewsGPT4 84.95 73.10 74.81 77.62
Table 5: Ablation study of prompt strategies. AUC (in %) of variant models is reported on the remaining three subsets when trained on the BBC subset. SPT denotes soft prompt tuning, whereas CAH refers to candidate answer heuristics.
4.3 Ablation Study
We perform several ablation experiments to explore the necessity of the proposed component modules, prompt strategies, and LVLMs knowledge, respectively. In the following experiments, all ablation results are evaluated on the remaining three subsets when trained on the BBC subset.
The Effect of the Component Modules.
In Table 4, we conduct a comprehensive ablation study on the proposed component modules to verify their effectiveness. The first row of Table 4 shows our baseline model that only performs soft prompt tuning, achieving an average AUC of 52.5%. Based on this baseline, we further introduce two separate modules: a multi-level cross-modal (ML&CM) reasoning module and a dual-branch fine-grained verification (DB&FG) module, with 21.55% and 20.16% improvement in average AUC, respectively. This implies that forgery-specific knowledge augmentation is indispensable for our framework. Comparatively, our model with complete two modules obtains the best performance increasing by 25.12%, indicating the effectiveness and complementarity of these two modules. Moreover, we test the performance of FakeNewsGPT4 removing the multi-level features (w/o ML), and FakeNewsGPT4 removing any one of the dual-branch features (w/o DB). These variant models lack rich features to represent forgery-specific knowledge leading to a great decrease in cross-domain performance.
The Effect of the Prompt Strategies.
The prompt strategies designed in the alignment process and the corresponding results for each case are tabulated in Table 5. First, when removing continuous prompt vectors (w/o SPT), the performance drops a little bit. In addition, after removing the candidate answer options (w/o CAH) in the human instruction prompt, the average performance decreases from 77.62% to 73.69%. In particular, the third row of Table 5 represents that none of both strategies is employed in our proposed framework. Our method substantially outperforms this variant model, implying both two strategies enable the introduction of implicit information to fully activate the capacity of LVLMs.
Figure 4: Ablation study of the world knowledge inherent in large vision-language models.
The Effect of the LVLMs Knowledge.
To analyze the impact of world knowledge derived from LVLMs, we compare our method with the common practice of processing fused embeddings in a supervised classification manner Shao et al. (2023a). This variant model (w/o World Knowledge) replaces the Vicuna model in FakeNewsGPT4 with a binary classifier to predict true/fake labels. As shown in Fig. 4, harnessing the inherent knowledge in LVLMs improves an average performance by 9.92 points over the variant model. Furthermore, for target domains from distinct countries exhibiting huge distribution differences, FakeNewsGPT4 yields more significant improvements in 13.26% and 12.15%. This could be attributed to the fact that world knowledge from LVLMs can effectively guide the representation of agnostic instances.
5 Conclusion
Our work presents FakeNewsGPT4, a novel framework that leverages world knowledge from LVLMs and forgery-specific knowledge augmentation to alleviate the domain shift issue in multimodal fake news detection. Two types of critical knowledge for manipulation reasoning are augmented in FakeNewsGPT4. Specifically, the multi-level cross-modal reasoning module is devised to focus on the cross-modal semantic correlations, while the dual-branch fine-grained verification module gathers local spatial information to encode artifact traces. Then, we transform the generated knowledge into refined embeddings for alignment with language space. The candidate answer heuristics and soft prompts are introduced as supplementary inputs to unleash the extensive knowledge of LVLMs. Extensive experiments verify that FakeNewsGPT4 shows superior cross-domain performance compared to the state-of-the-art methods.
Limitations
While the proposed FakeNewsGPT4 framework demonstrates promising results in multimodal fake news detection, there are several limitations to consider. Firstly, we capitalize on the inherent world knowledge derived from large vision-language models (LVLMs), which learn from extensive internet data, to alleviate the domain shift issue. Since the open world is constantly evolving, we argue that the cross-domain performance may be affected by the lack of updates to world knowledge in LVLMs when confronted with future fake news. In the future, we can try to update world knowledge of LVLMs by model editing techniques. Secondly, our current approach only considers two modalities, vision and text, in multimodal fake news detection tasks. The malicious manipulation of other modalities such as audio, also poses a significant threat to people’s daily lives. Future studies can extend our experiments to explore the joint detection of AIGC data in more modal combinations.
Ethics Statement
It is crucial to clarify that the categorization of fake news relies on the source information provided in the VisualNews dataset Liu et al. (2021a). VisualNews dataset consists of image-text pairs collected from multiple online news agencies and is publicly available for research purposes only. All experiments are performed on publicly available datasets which are commonly used in research on multimodal fake news detection. Hence, we do not think of any potential harm or bias that can be derived from our work.
Appendix
Appendix A Contrastive Class Prompts
Fig. 5 presents a detailed list of prompts utilized for implementing the contrastive class prompt which is introduced in Section 3.1.2 of the main text. Following WinCLIP Jeong et al. (2023), we employ the compositional prompt ensemble to generate texts representing natural and unnatural states. Specifically, we curate prompt templates involving a photo of a [c] and a photo of the [c]. A complete prompt can be composed by replacing the token [c] in the template-level prompt with one of the state-level prompts, either from the natural or unnatural states.
Figure 5: Lists of two state-level prompts considered in this paper to construct contrastive class prompts.
Appendix B Verification Loss
Pixel-wise Verification Loss
The manipulated segmentation map $M_{s}$ is utilized to calculate focal loss Lin et al. (2017) and dice loss Milletari et al. (2016) supervised by the grounding mask. In the multimodal fake news detection task, where the majority of regions in fake images remain pristine, employing focal loss can alleviate the issue of class imbalance. The Focal loss is computed as follows:
$\mathcal{L}_{\text{focal}}=-\frac{1}{n}\sum_{i=1}^{n}{(1-p_{i})^{\gamma}{\rm log}(p_{i})},$ (10)
where $n$ denotes the total number of pixels, $p_{i}$ represents the predicted probability of positive classes, and $\gamma$ is a hyperparameter for adjusting the weight of hard-to-classify samples. In our implementation, we set $\gamma$ to 2.
Dice loss is based on the dice coefficient and can be computed as follows:
$\mathcal{L}_{\text{dice}}=1-\frac{2\sum_{i=1}^{n}{y_{i}\hat{y}_{i}}}{\sum_{i=1}^{n}{y_{i}^{2}}+\sum_{i=1}^{n}{\hat{y}^{2}_{i}}},$ (11)
where $n$ denotes the total number of pixels, $y_{i}$ is the pixel value in the segmentation map and $\hat{y}_{i}$ is the ground truth value.
Path-level Verification Loss
To regress the predicted bounding box, the aggregated token $u_{{\rm agg}}$ is fed into the BBox Detector $D_{v}$, which comprises two multi-layer perception (MLP) layers. Then we compute the patch-level verification loss by combining normal L1 loss and generalized Intersection over Union (IoU) loss as follows:
$\hat{b}$ $=D_{v}\left(u_{\rm agg}\right),$ (12)
$\mathcal{L}_{\text{patch}}$ $=\mathcal{L}_{\mathrm{L}_{1}}(b,\hat{b})+\mathcal{L}_{\text{giou}}(b,\hat{b}),$
where $\hat{b}$ denote the predicted bounding boxes and $b$ denote the ground-truth box.
Appendix C Real-world Distribution Divergence
In Fig. 6, we present word clouds to illustrate the distribution divergence in real-world context across regional perspectives and thematic focus. Notably, the BBC predominantly reports British news, encompassing various subjects such as culture and entertainment. Conversely, the Washington Post tends to emphasize American political news. This disparity leads to variations in word usage across distinct domains. For instance, the commonly used words in BBC news include “prime minister", “children", and "family" etc, while in the Washington Post news, prevalent terms are “president", “Donald Trump", and “Obama", etc.
Figure 6: Word clouds of training data sourced from subsets of BBC, The Guardian, USA Today, and The Washington Post, where the size of terms corresponds to the word frequency.
Appendix D Implementation Details
Here we provide detailed implementation details for training on the remaining three subsets: The Guardian, USA TODAY, and The Washington Post. The hyperparameters are the same for all subsets but differentiate the learning rate, batch size, and maximum epoch by dataset scale, detailed in Table 6. Additionally, to ensure a fair comparison, the hyperparameters for the PandaGPT using soft prompt tuning are set to be consistent with the values used in our method. For the state-of-the-art method HAMMER, the hyperparameters are set to default values used in  Shao et al. (2023a).
Parameters Train set
Guardian USA Wash.
Optimizer AdamW AdamW AdamW
Learning Rate 4.1e-5 1.1e-5 1e-5
LR Scheduler Linear Linear Linear
Batch Size 128 16 16
Maximum Epoch 12 10 10
Table 6: Hyperparameters used in The Guardian, USA TODAY, and The Washington Post subsets.
Appendix E Case Analysis
Fig. 7 and Fig. 8 depict cases from the testing set. The former illustrates comparisons between the vanilla LVLM and FakeNewsGPT4, while the latter showcases cases where either the compared methods or the LVLM using soft prompt tuning predictions are incorrect.
In Fig. 7, Case 1 and Case 2 show cases where the vanilla LVLM model predicts all labels as “Fake” while the FakeNewsGPT4 makes correct predictions. This implies that the vanilla LVLM lacks the necessary knowledge to effectively detect fake news when learning from the general corpus, making it challenging to accurately characterize the concept of forgery and provide precise responses. Consequently, the vanilla LVLM relies heavily on common prompt words, such as “face”, resulting in all labels being predicted as “Fake”.
Figure 7: Cases in the testing set where vanilla LVLM and FakeNewsGPT4 confront with pristine news and fake news. ( ) indicates wrong prediction and ( ) indicates correct prediction.
Figure 8 illustrates comparisons between our method with the state-of-the-art method HAMMER, and PandaGPT using soft prompt tuning (PandaGPT+SPT). In Case 3, the absence of forgery-specific knowledge impedes the ability of PandaGPT to perform manipulation reasoning, especially when confronted with fake images containing subtle artifacts. In Case 4, the powerful model HAMMER struggles to handle agnostic instances due to the lack of prior information, resulting in incorrect judgments regarding unseen fake news. Conversely, Case 5 demonstrates the effectiveness of our FakeNewsGPT4, which integrates inherent world knowledge from LVLMs and incorporates forgery-specific information to make accurate predictions.
Figure 8: Cases in the testing set where at least one in the Baseline and the LVLM using soft prompt tuning made incorrect predictions. ( ) indicates wrong prediction and ( ) indicates correct prediction.